Solve an equilibrium problem (using an ICE table) to calculate the pH of of each solution:

a solution that is 0.14 M in NH3 and 0.19 M in NH4Cl
Answer #1

The equilibrium expression for the dissociation of ammonium chloride is shown below as:-

NH_4Cl\rightleftharpoons NH_3+HCl

Consideirng the ICE table as:

\begin{matrix} & NH_4^+ & \rightleftharpoons & NH_3 & + & H^+\\ Initial& 0.19\ M & &0.14\ M & & \\ Change\ in\ concentration &-x & & +x & & +x\\ Equilibrium\ concentration& 0.19\ M-x& &0.14\ M +x & & +x \end{matrix}

The expression for Ka is shown below as:-

K_a=\frac{[NH_3][H^+]}{[NH_4^+]}

The value of Ka of ammonium chloride is 5.6\times 10^{-10} .

5.6\times 10^{-10}=\frac{x\times (0.14+x)}{(0.19-x)}

Solving for x, we get that:-

x = 7.6\times 10^{-10} .

pH=-\log[H^+]=-\log \left(7.6\times 10^{\left\{-10\right\}}\right)=9.12